The resistance of a coil of a copper wire at the beginning of a heat test is 173Ω, the temperature being 160C. At end of the test, the resistance is 212Ω. Calculate the temperature raise of the coil. Assuming the temperature coefficient of resistance of copper to be 0.00426/0C at 00C.


1
Expert's answer
2021-09-02T10:42:34-0400

At 16°C and another temperature:


"R_1=R_0(1+\alpha T_1),\\
R_2=R_0(1+\alpha T_2)."

Express T2:


"T_2=\frac 1\alpha\bigg(\frac{R_2}{R_0}-1\bigg),"

from the first expression we have


"R_0=\frac{R_1}{1+\alpha T_1}.\\\space\\
T_2=\frac 1\alpha\bigg(\frac{R_2}{\frac{R_1}{1+\alpha T_1}}-1\bigg)=73°\text{C}."